A square quilt has a border that is 1 foot wide on each side. The quilt has an area of 25 square feet. What is the length of the quilt without the border?
Answer by ankor@dixie-net.com(22740)

A square quilt has a border that is 1 foot wide on each side.
The quilt has an area of 25 square feet.
What is the length of the quilt without the border?
:
Let x = length of each side without the border
:
A 1' border adds 2' to each dimension of the inner square
:
(x+2)^2 = 25
FOIL
x^2 + 4x + 4 = 25
x^2 + 4x + 4 - 25 = 0
x^2 + 4x - 21 = 0
Factors to
(x+7)(x-3) = 0
positive solution is what we want here
x = 3 ft on each side without the border